OpenStudy (anonymous):

factor: x^2 - 3x - 10 factor: x^2 -16

jhonyy9 (jhonyy9):

x^2 -3x -10 =(x-5)(x+2)

jhonyy9 (jhonyy9):

for second us formula 2 -b^2 =(a-b)(a+b) ok ?

jhonyy9 (jhonyy9):

oh sorry have list the a

jhonyy9 (jhonyy9):

so a^2 -b^2 =(a-b)(a+b)

jhonyy9 (jhonyy9):

ok ?

jhonyy9 (jhonyy9):

this is easy a=x b=4

jhonyy9 (jhonyy9):

try it now alone